OpenStudy (mwhit2101):

SEVENTH GRADE Inequality Help Medal+Fan for best answer Write an inequality to represent each statement. 1. A number less than n incrased by 6 is less than or equal to 4.

OpenStudy (anonymous):

hdehdhdhdhdhhldhjlsdv

OpenStudy (mwhit2101):

Great to know, thank you

OpenStudy (rational):

confusing :S

OpenStudy (mwhit2101):

How is it cofusing

OpenStudy (rational):

1. A number less than n incrased by 6 is `less than or equal to 4.` that part translates to \(\large \le 4\)

OpenStudy (mwhit2101):

I just don't understand the "a number less than n" I understand the rest though. If there was no "a number less than n" it would simply be \[n+6\]

OpenStudy (mwhit2101):

oops

OpenStudy (mwhit2101):

\[n + 6 \le 4\]

OpenStudy (mwhit2101):

But i dont understand the "a number less than"

OpenStudy (anonymous):

1. A number less than n incrased by 6 is less than or equal to 4. \[N-x + 6 \le 4 \] ....